OpenStudy (anonymous):

describe how to transform (^3 sqrt x^4)^5 into an expression with a rational exponent. make sure you respond with complete sentences

OpenStudy (anonymous):

Still need help? Note: it might be better if you try to write your expression with the Equation writer.

OpenStudy (anonymous):

\[\left( \sqrt[3]{X ^{4}} \right)^{5}\]

OpenStudy (anonymous):

this is the equation please help

OpenStudy (anonymous):

Which part confuses you?

OpenStudy (anonymous):

i wouldn't know how to transform it into an expression with a rational exponent.

OpenStudy (anonymous):

first, do you know how to write \[\sqrt[3]{x} or \sqrt{x} \] as an exponent?

OpenStudy (anonymous):

First, if you have something like: \[y = \sqrt[3]{x}\] that means that: \[y^{3} = x\] Are you with me so far?

OpenStudy (anonymous):

yes so far i was kind of behind with my class so anything helps please

OpenStudy (anonymous):

Also, do you know how to simplify an expression like this: \[(x^{3})^{4}\]

OpenStudy (anonymous):

nope please explain

OpenStudy (anonymous):

what about: \[x^{3}*x^{3}\] Do you know what this equals?

OpenStudy (anonymous):

yes \[X^{3}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Not quite...

OpenStudy (anonymous):

Here's an example: What's 8*8?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

Because \[8 = 2^{3}\] And 8*8 is the same as: \[2^{3}*2^{3}\]

OpenStudy (anonymous):

So I'll give you a second shot: \[x^3*x^3 = ?\]

OpenStudy (anonymous):

\[X^{6}\]

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

took me a min little slow

OpenStudy (anonymous):

Another way of saying that is that: \[x^a*x^b = x^{a+b}\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

And, with exponents, \[(x^a)^3 = x^a*x^a*x^a = x^{a+a+a} = x^{3a}\]

OpenStudy (anonymous):

So, \[x^a*x^b = x^{a+b}\] BUT, \[(x^a)^b = x^{ab}\]

OpenStudy (anonymous):

Are you getting a hang of this?

OpenStudy (anonymous):

now that your taking your time to explain it I'm starting to understand it

OpenStudy (anonymous):

Ok, great!

OpenStudy (anonymous):

Now, the hardest part about that first expression you gave me was the "cubed root". You have any idea how to deal with those??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

First, I want you to simplify these two expressions while I type something about cubed roots: \[x^3*2x^2\] and \[(x^4)^7\]

OpenStudy (anonymous):

\[2x^{5}\]

OpenStudy (anonymous):

First, lets say I give you something like: \[(x^a)^3\] a.) first, simplify this b.) if I say that: \[(x^a)^3= x\] what do you think "a" equals.

OpenStudy (anonymous):

\[\left(x ^{47} \right)\]

OpenStudy (anonymous):

a equals 4

OpenStudy (anonymous):

Whoops, not quite. You got the first question right (I even threw in that extra "2" to trip you up).

OpenStudy (anonymous):

Look back at that second question -- (x^4)^7

OpenStudy (anonymous):

\[\left( X^{27} \right)\]

OpenStudy (anonymous):

You're almost there, except remember: the exponent equals 4*7,

OpenStudy (anonymous):

i mean 28 sorry

OpenStudy (anonymous):

Ok, great!

OpenStudy (anonymous):

Here's another one: If \[x^a = x\] Then what is "a" equal to?

OpenStudy (anonymous):

\[x^{2}\]

OpenStudy (anonymous):

So you're saying: \[x^2 = x\] ?

OpenStudy (anonymous):

i believe so.

OpenStudy (anonymous):

Now, that IS true when x = 0 or when x =1.... but what about for ALL x's

OpenStudy (anonymous):

Well, try plugging in a number for "x" to see. For instance, try out "x=3"

OpenStudy (anonymous):

will it be 9?

OpenStudy (anonymous):

You're right: 3^2 = 9. Now, let's go back to that original question: \[x^a = x\] If "a" was equal to "2" then that would mean that: \[x^2 = x\] And, if we plugged in "3" for "x" then you would have: \[3^2 = 3\] But you just said that 3^2 = 9 (which is true). Therefore, \[3^2 \neq 3\]

OpenStudy (anonymous):

This might sound boring but I swear this part is important for when you turn the "cubed root" into an exponent...

OpenStudy (anonymous):

sorry was in restroom

OpenStudy (anonymous):

Next time you can just say "I was busy". :)

OpenStudy (anonymous):

sorry but thank you

OpenStudy (anonymous):

I'll give you a second to catch up -- I have to take some medicine myself (I'm sick).

OpenStudy (anonymous):

okay thank you and sorry to hear that i just got over a cold myself.

OpenStudy (anonymous):

Ok, now back to that question I gave you earlier: \[x^a = x\]

OpenStudy (anonymous):

\[x^{a2}\]

OpenStudy (anonymous):

Not quite. But the good news is that it's easier than it looks.

OpenStudy (anonymous):

a = 1 Because \[x^1 \] is just another way of writing "x"

OpenStudy (anonymous):

ohh yea duh i should of got that right

OpenStudy (anonymous):

Because if: \[x^3 = x*x*x\] \[x^2 = x*x\] Then \[x^1 = x\]

OpenStudy (anonymous):

Now, back to roots, what is: \[\sqrt[3]{x}*\sqrt[3]{x}*\sqrt[3]{x}\]

OpenStudy (anonymous):

\[\sqrt[27]{x}\]

OpenStudy (anonymous):

\[\sqrt[27]{X}^{3}\]

OpenStudy (anonymous):

How about this one: \[\sqrt{x}*\sqrt{x}\]

OpenStudy (anonymous):

Hint: It's easier than it looks...

OpenStudy (anonymous):

And it always helps if you plug in some numbers: \[\sqrt{16}*\sqrt{16} = ?\]

OpenStudy (anonymous):

but how do you always know which numbers to plug in?

OpenStudy (anonymous):

No, whenever you see an "x" that means that it could stand for ANY number. Thus, I can choose any number I want just to test my answers. There's nothing special about the "16" I chose--I chose it because the square root of 16 is easy to calculate (it's just "4"). Similarly, when talking about cube roots, I would probably chose an a number like "27" because I know from the top of my head that the cube root of 27 equals 3.

OpenStudy (anonymous):

Now, if I gave you something like: \[9 = \sqrt{x}\] You can't just plug in any number you want for "x" --you have to SOLVE for x. But, if you just have regular expression (i.e. with no "equals" sign) then you can always plug in a number for "x" and see if your solution makes sense.

OpenStudy (anonymous):

Remember how I earlier wrote that: \[y = \sqrt[3]{x}\] means that: \[y ^{3} = x\]? Well, that means that: \[\sqrt[3]{y^3} = x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually, sometimes it's most helpful to work backwards. Because it will all make sense after I give you the answer... First, to rewrite a cubed root as an exponent: \[\sqrt[3]{x} = x^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

and, because: \[(x^a)^b = x^{a*b}\] That means that you can take any type of root such as: \[\sqrt[7]{x^5}\] and rewrite it as: \[(x^5)^\frac{ 1 }{ 7 }\] which than can be simplified to: \[x^\frac{ 5 }{ 7 }\]

OpenStudy (anonymous):

So to go back to your original question: \[(\sqrt[3]{x^4})^5 = ((x^4)^\frac{ 1 }{ 3 })^5 = (x^\frac{ 4 }{ 3 })^5 = x^\frac{ 20 }{ 3 }\] Thus your answer is "x raised to the twenty thirds power" or similarly, "x to the 6 and two thirds power"

OpenStudy (anonymous):

i truly thank you so much you are the biggest help I've ever had truly better than my own teachers but then again I'm in online so i guess I'm my own teacher and i suck lol

OpenStudy (anonymous):

Well, exponents can be confusing... Good luck with the rest of your work!

OpenStudy (anonymous):

thanks